OpenStudy (anonymous):

give the next three numbers in the pattern 1,3,7,15,2,4,6

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

wrong !!!

OpenStudy (anonymous):

Is the pattern 1, 3, 7, 15, ---------- ??

OpenStudy (anonymous):

1 + 2^1 = 1 + 2 = 3 3 + 2^2 = 3 + 4 = 7 7+ 2^3 = 7 + 8 = 15 15 + 2^4 = 15 + 16 = 31 31 + 2^5 = 31 + 32 = 63 63 + 2^6 = 63 + 64 = 127